# Unraveling particle dark matter with Physics-Informed Neural Networks

### v1

@@ 5.2 Discovering power-law cosmologies
Figure 9: PINN solution for the inverse problem (see text for details): relation between $\gamma$ and $C$, for $z_{t}=\ln 10$ and switch transition (16) in BE (18), satisfying $\Omega h^{2}=0.12$ [3]. The dashed-line was obtained via linear interpolation of the points. Colored point in green (blue) [red] indicate the case of GB (Standard) [RS] cosmology (see Figs. 7 and 8), while the black point values are in Table 2.
$C$ $\langle\sigma v\rangle$ $\left(\text{GeV}^{-2}\right)$ $\gamma_{\text{PINN}}$ $W_{n}(z_{f})_{\text{FEM}}$ $\Omega h^{2}_{\text{FEM}}$
@@ 5.3 Modeling alternative cosmology: transition between power and Hubble law
Hence, $\mathrm{Softplus}_{\beta}(x)$ is a smooth approximation to ReLU, where the “speed" of the transition is controlled by $\beta$, recovering the switch transition in the limit $\beta\rightarrow\infty$. The middle point of this transition is fixed at $x=0$, and is independent of the choice of $\beta$. Thus, we can define a new prameterization function – $G(x,x_{t},\gamma)$ – that models the transition between power and Hubble law in alternative cosmology:
@@
$\gamma$ $C_{\text{Switch}}$ $\langle\sigma v\rangle$ (×$\left(\times 10^{-27}\ \text{GeV}^{-2}\right)$ $W_{n}(z_{f})_{\text{FEM}}$ $\Omega h^{2}_{\text{FEM}}$
@@
$\gamma$ $C_{\text{Smooth}}$ $\langle\sigma v\rangle$ (×$\left(\times 10^{-27}\ \text{GeV}^{-2}\right)$ $W_{n}(z_{f})_{\text{FEM}}$ $\Omega h^{2}_{\text{FEM}}$
@@ 5.4 Bayesian PINN: a method to quantify epistemic uncertainties
Although this uncertainty is parametric, i.e. parametrized by $\lambda_{\text{KL}}$, we found it to be useful in cases where FEMs are not entirely trustworthy. It is also noteworthy that without the KL divergence, the posterior was also found not to collapse entirely. The remaining small amount of uncertainty is then completely related to the optimization process, and was found to be of the order of $10^{-4}$, expected given our previous results in the non-Bayesian inverse PINN [see Section 5.1].
@@ 6 Conclusions and outlook
In conclusion, PINNs provide a valuable model-building theoretical tool useful in unraveling theories that can explain experimental data. Future directions include the full resolution of coupled systems, such as modified Friedman equations in the context of alternative cosmologies and BE governing DM abundance. Moving towards a complete particle physics model, where we compute the interactions cross-sections, could ultimately make PINNs an interesting alternative to the conventional numerical packages for DM studies. The inclusion of multiple experimental constraints, besides the CDM relic abundance, in a concrete model realization would be a challenging but interesting problem for inverse PINNs. Possibilities for further applications of this algorithm in the realms of cosmology, particle and astroparticle physics are endless.

### v2

@@ 5.2 Discovering power-law cosmologies
Figure 10: PINN solution for the inverse problem (see text for details): relation between $\gamma$ and $C$, for $z_{t}=\ln 10$ and switch transition (16) in BE (18), satisfying $\Omega h^{2}=0.12$ [3]. The dashed-line was obtained via linear interpolation of the points. Colored point in green (blue) [red] indicate the case of GB (Standard) [RS] cosmology (see Figs. 8 and 9), while the black point values are in Table 2.
$C$ $\langle\sigma v\rangle$ $\left(\text{GeV}^{-2}\right)$ $\gamma_{\text{PINN}}$ $W(z_{f})_{\text{FEM}}$ $\Omega h^{2}_{\text{FEM}}$
@@ 5.3 Modeling alternative cosmology: transition between power and Hubble law
Hence, $\mathrm{Softplus}_{\beta}(x)$ is a smooth approximation to ReLU, where the “speed" of the transition is controlled by $\beta$, recovering the switch transition in the limit $\beta\rightarrow\infty$. The middle point of this transition is fixed at $x=0$, and is independent of the choice of $\beta$. Thus, we can define a new parameterization function – $G(x,x_{t},\gamma)$ – that models the transition between power and Hubble law in alternative cosmology:
@@
$\gamma$ $C_{\text{Switch}}$ $\langle\sigma v\rangle$ (×$\left(\times 10^{-27}\ \text{GeV}^{-2}\right)$ $W(z_{f})_{\text{FEM}}$ $\Omega h^{2}_{\text{FEM}}$
@@
$\gamma$ $C_{\text{Smooth}}$ $\langle\sigma v\rangle$ (×$\left(\times 10^{-27}\ \text{GeV}^{-2}\right)$ $W(z_{f})_{\text{FEM}}$ $\Omega h^{2}_{\text{FEM}}$
@@ 5.4 Bayesian PINN: a method to quantify epistemic uncertainties
Although this uncertainty is parametric, i.e. parameterized by $\lambda_{\text{KL}}$, we found it to be useful in cases where FEMs are not entirely trustworthy. It is also noteworthy that without the KL divergence, the posterior was also found not to collapse entirely. The remaining small amount of uncertainty is then completely related to the optimization process, and was found to be of the order of $10^{-4}$, expected given our previous results in the non-Bayesian inverse PINN [see Section 5.1].
@@ 6 Conclusions and outlook
Although inverse PINNs offer a flexible and powerful alternative to traditional FEMs, several limitations remain. The performance of PINNs is highly sensitive to the choice of hyperparameters, that require extensive manual tuning in the absence of a systematic optimization strategy. Furthermore, handling stiffness in differential equations remains a significant challenge. For instance, while the freeze-in BE considered here is not explicitly stiff, the freeze-out regime leads to Riccati-type equations that are explicitly stiff, with sharp transitions in the solution requiring careful numerical treatment. PINN architectures also need to be tailored to the specific form of the underlying ODEs or PDEs, limiting generalizability. Additionally, training PINNs can be computationally expensive, particularly when high accuracy is required or when dealing with complex systems involving multiple coupled equations. Addressing these limitations is essential for scaling PINNs to broader applications in particle physics and cosmology.
In conclusion, PINNs provide a valuable model-building theoretical tool useful in unraveling theories that can explain experimental data. Future directions include the full resolution of coupled systems, such as modified Friedman equations in the context of alternative cosmologies and BE governing DM abundance. Moving towards a complete particle physics model, where we compute the interactions cross-sections, could ultimately make PINNs an interesting alternative to the conventional numerical packages for DM studies. A promising avenue is the development of modular PINNs [117], in which multiple networks are linked – for example, one network could approximate the thermally averaged cross section, while another solves the BE to extract the particle abundance. Furthermore, the inclusion of multiple experimental constraints, besides the CDM relic abundance, in a concrete model realization would be a challenging but interesting problem for inverse PINNs. Possibilities for further applications of this algorithm in the realms of cosmology, particle and astroparticle physics are endless.